Cos(a+b)=0.8 Sin(a-b)5/13 and find Tan2a

Grade:10

2 Answers

Sunil Raikwar
9 years ago
sin2a=sin{(a+b)+(a-b)}=sin(a+b)cos(a-b)+cos(a+b)sin(a-b)
=3/5.12/13+4/5.5/13
=56/65
tan2a=56/33
Thanks & Regards
Sunil Raikwar
askIITians Faculty
siddharth gupta
9 years ago
cos(a+b)=4/5i.e cos-14/5=a+b i.e tan-13/4=a+b(can be calculated by considering a right triangle)......(1)
similarly sin-15/13=a-b and hence tan-15/12=a-b(calculated as above)......(2)
add (1)and(2)
a+b+a-b=2a=tan-15/13+tan-15/12=tan-156/33i.e tan2a=56/33.
thanks
siddharth gupta